The value of 2-2*3/5 is


1
Expert's answer
2021-05-07T10:09:24-0400

Using BODMAS Rule:




Divide and Multiply rank equally (and go left to right).

Add and Subtract rank equally (and go left to right)


So, "2-2\times 3/5"


Step1: First solve multiplication according to rule

"2-2\times 3/5=2-6/5"


Step 2: Now Divide

"2-6\/5=2-1.2"


Step3: Then finally subtract

2"-1.2=0.8"


Final result:

2-2X3/5=0.8